# AI Research Agents for Machine Learning: Search, Exploration, and Generalization in MLE-bench

### v1

@@ 6 Related Works
Scaling Search with LLMs. Integrating search with LLM-based generators has gained traction across domains such as coding (Antoniades et al., 2025), math (Yao et al., 2023a), and planning (Hazra et al., 2024; Hao et al., 2023). In practice, increased test-time compute through algorithms like best-of-$N$(Wang et al., 2023), beam search (Hazra et al., 2024), MCTS (Antoniades et al., 2025; Hao et al., 2023), and evolutionary search (Romera-Paredes et al., 2024; Hazra et al., 2025) often improves performance beyond the architectural and size constraints of LLMs (Snell et al., 2024). Our results highlight that these performance gains depend critically on the interplay between search components (see Section 8)—an aspect that prior work has largely overlooked.
@@
Automating ML Engineering and Scientific Discovery. Traditional approaches like AutoML (Feurer et al., 2022; Thornton et al., 2013; Olson and Moore, 2019) and Neural Architecture Search (NAS)(Zoph and Le, 2017; Liu et al., 2019; Real et al., 2019) automate ML by searching over predefined, expert-designed configuration spaces, often via brute-force or heuristic methods (Bergstra and Bengio, 2012; Elsken et al., 2017; Li et al., 2018). In contrast, recent advances in LLMs enable more open-ended design. AIDE (Jiang et al., 2025) treats ML engineering as LLM-guided code optimization via tree search. AI Scientist v2 (Yamada et al., 2025) extends this paradigm, automating the entire research pipeline using agentic LLMs. Similar techniques have been applied to software engineering (Antoniades et al., 2025), algorithm and reward design (Romera-Paredes et al., 2024; Lu et al., 2024; Faldor et al., 2025; Hazra et al., 2025; DeepSeek-AI, 2025), and even natural sciences (Boiko et al., 2023; Swanson et al., 2024). Existing systems often combine search procedures, operators, and evaluation in ways that make it challenging to understand which components drive performance improvements. To address this, we separate these components and look at how each one—and their interactions—can be optimized better. Concurrent work, R&D-Agent (Yang et al., 2025), which addresses the same problem setting, achieved impressive results on MLE-Bench. Our results fall within their reported standard deviation. We note that their experiments use at most 6 seeds, and as discussed in Section 15, the limited number of seeds may introduce variation in the conclusions. Finally, our approach differs from methods such as Agent K (Grosnit et al., 2024), which uses long-term memory across multiple Kaggle competitions. In contrast, we focus on addressing each competition independently.
@@
AI Research Frameworks. Most existing benchmarks for machine learning or research engineering, such as MLGym (Nathani et al., 2025), RE-bench (Wijk et al., 2024), and MLAgentBench (Huang et al., 2023), come with their own frameworks. Our approach is most similar to Inspect (Institute, ), a framework which provides an abstraction that makes minimal assumptions about the agent and evaluation design, clearly separating the two. This flexibility in agent design is essential for enabling rigorous comparisons across a wide range of agents—from simple LLM-based agents with tool access to scaffolds combining algorithmic and LLM-based components—within the same environment. The environment plays a critical role in performance (see Section 5.2). However, unlike Inspect, AIRA-dojo focuses strongly on long-running research tasks, which influences our environment design choices. For example, prior work typically uses Docker to containerize agents’ workspaces, but Docker is unsupported on most HPC clusters due to its reliance on root privileges and lack of seamless integration with common HPC resource managers. This limitation hinders scalability, which we address through our Apptainer-based solution (see Section 3.2).
@@ 7.1 Limitations and Future Work
Scaling the search. To maintain comparability with MLE-bench, we adopt the benchmark’s time (24-hour limit) and compute constraints (1 GPU). However, solving challenging problems likely requires substantially greater resources, and evaluating performance under these restrictions provides limited insights (see Section 10 for an experiment extended to 120 hours—5 days). Studying the scaling behavior of search policies and operators, and developing agents capable of effectively leveraging more computational resources over longer time horizons, is an important direction for future research.

### v2

@@ 6 Related Works
Scaling Search with LLMs. Integrating search with LLM-based generators has gained traction across domains such as coding (Antoniades et al., 2025), math (Yao et al., 2023a), and planning (Hazra et al., 2024; Hao et al., 2023). In practice, increased test-time compute through algorithms like best-of-NN(Wang et al., 2023), beam search (Hazra et al., 2024), MCTS (Antoniades et al., 2025; Hao et al., 2023), and evolutionary search (Romera-Paredes et al., 2024; Hazra et al., 2025) often improves performance beyond the architectural and size constraints of LLMs (Snell et al., 2024). Our results highlight that these performance gains depend critically on the interplay between search components (see Section˜8)—an aspect that prior work has largely overlooked.
@@
Automating ML Engineering and Scientific Discovery. Traditional approaches like AutoML (Feurer et al., 2022; Thornton et al., 2013; Olson and Moore, 2019) and Neural Architecture Search (NAS)(Zoph and Le, 2017; Liu et al., 2019; Real et al., 2019) automate ML by searching over predefined, expert-designed configuration spaces, often via brute-force or heuristic methods (Bergstra and Bengio, 2012; Elsken et al., 2017; Li et al., 2018). In contrast, recent advances in LLMs enable more open-ended design. AIDE (Jiang et al., 2025) treats ML engineering as LLM-guided code optimization via tree search. AI Scientist v2 (Yamada et al., 2025) extends this paradigm, automating the entire research pipeline using agentic LLMs. Similar techniques have been applied to software engineering (Antoniades et al., 2025), algorithm and reward design (Romera-Paredes et al., 2024; Lu et al., 2024; Faldor et al., 2025; Hazra et al., 2025; DeepSeek-AI, 2025), and even natural sciences (Boiko et al., 2023; Swanson et al., 2024). Existing systems often combine search procedures, operators, and evaluation in ways that make it challenging to understand which components drive performance improvements. To address this, we separate these components and look at how each one—and their interactions—can be optimized better. Concurrent work, R&D-Agent (Yang et al., 2025), which addresses the same problem setting, achieved impressive results on MLE-Bench. Our results fall within their reported standard deviation. We note that their experiments use at most 6 seeds, and as discussed in Section˜16, the limited number of seeds may introduce variation in the conclusions. Finally, our approach differs from methods such as Agent K (Grosnit et al., 2024), which uses long-term memory across multiple Kaggle competitions. In contrast, we focus on addressing each competition independently.
@@
AI Research Frameworks. Most existing benchmarks for machine learning or research engineering, such as MLGym (Nathani et al., 2025), RE-bench (Wijk et al., 2024), and MLAgentBench (Huang et al., 2023), come with their own frameworks. Our approach is most similar to Inspect (Institute, 2025), a framework which provides an abstraction that makes minimal assumptions about the agent and evaluation design, clearly separating the two. This flexibility in agent design is essential for enabling rigorous comparisons across a wide range of agents—from simple LLM-based agents with tool access to scaffolds combining algorithmic and LLM-based components—within the same environment. The environment plays a critical role in performance (see Section˜5.2). However, unlike Inspect, AIRA-dojo focuses strongly on long-running research tasks, which influences our environment design choices. For example, prior work typically uses Docker to containerize agents’ workspaces, but Docker is unsupported on most HPC clusters due to its reliance on root privileges and lack of seamless integration with common HPC resource managers. This limitation hinders scalability, which we address through our Apptainer-based solution (see Section˜3.2).
@@ 7.1 Limitations and Future Work
Scaling the search. To maintain comparability with MLE-bench, we adopt the benchmark’s time (24-hour limit) and compute constraints (1 GPU). However, solving challenging problems likely requires substantially greater resources, and evaluating performance under these restrictions provides limited insights (see Section˜10 for an experiment extended to 120 hours—5 days). Studying the scaling behavior of search policies and operators, and developing agents capable of effectively leveraging more computational resources over longer time horizons, is an important direction for future research.